OpenStudy (anonymous):

Sophie has a rectangular garden with a length of (x - 5) and a width of (x + 4). Use the FOIL method to come up with the area of the garden. Your answer should be a simplified polynomial without parentheses. Explain how you used the FOIL method to get your answer.

OpenStudy (campbell_st):

just split it then distribute \[A = (x - 5)(x + 4)\] which is \[A=x(x + 4) - 5(x + 4)\] just distribute and collect like terms... its easier than foil

OpenStudy (anonymous):

\[A=lw\]\[l=(x-5)\]\[w=(x+4)\]\[A=(x-5)(x+4)\] \[FOIL=\text{First, Outside, Inside, Last}\]so you would have:\[(x\times x)(x\times4)(-5\times x)(-5\times4)\]which i think you should be able to complete :)